The cost of a pizza varies directly as the square of its radius. If a pizza with a 7 in radius costs $6.00, how much should a pizza with a 9-in radius cost?
Answer by stanbon(75887)

Directly related means Cost=k(radius^2)
To find "k" let cost be 6 and r be 7
Then 6=k(7^2)
So, k=6/49
The relating equation is therefore y=(6/49)r^2
So, if the radius is 9-inches the cost will be
C=(6/49)9^2
Cost=$9.92
Cheers,
stan H.